Sally was given 5 cents for each test she passed but had to return 10 cents for each she failed. At the end of 3 months she passed four times as many tests as she failed and she had earned 20 cents. How many tests did she pass?
Hi: My teacher handed this out today and is reprinted from a textbook, but I do not have the IBN number as the textbook is at school.
Thanks.

Answer by checkley77(12844)

5(4x)-10x=20
20x-10x=20
10x=20
x=20/10
x=2 failed tests.
4*2=8 passed tests.
Proof:
5*8-10*2=20
40-20=20
20=20